An object vibrates at the end of a vertical spring ( k = 23N/m). The speed of the object is 0.12m/s when the displacement is 0.15m. If the maximum displacement is 0.20m, what is the maximum speed of the object?
1
Expert's answer
2020-01-08T09:54:17-0500

Total energy of the system is "E = \frac{k a^2}{2}", where "k" is the spring constant, and "a" is the amplitude. This quantity is conserved, and is equal to the sum of kinetic and potential energies at any moment: "E = \frac{m v^2}{2} + \frac{k x^2}{2}".

For given displacement and velocity, the equation "E = \frac{m v_1^2}{2} + \frac{k x_1^2}{2}" holds.

The object reaches its maximum speed when "x = 0", then "E = \frac{m v_{max}^2}{2}", from where "v_{max} = \sqrt \frac{2 E}{m}".

Since we know the total energy in terms of amplitude "E = \frac{k a^2}{2}" , let us express the mass from the equation of energy for given position and velocity "m = \frac{ 2 E - k x_1^2}{v_1^2}"

, and substitute it into the expression for maximum velocity: "v_{max} = v_1 \sqrt{\frac{2 E}{2 E - k x_1^2}}" . Expressing energy in terms of the amplitude, and simplifying, obtain "v_{max} = v_1 \sqrt{\frac{a^2}{a^2 - x_1^2}}" .

Substituting "v_1 = 0.12 \frac{m}{s}", "a = 0.2 m", "x_1 = 0.15 m", and calculating, obtain "v_{max} \approx 0.18 \frac{m}{s}".